Faculty of Technical Sciences Maribor (TFM)

Department of Mechanical Engineering Maribor Slovenia

Researchers Anton Jezernik, Bojan Dolšak.

ILP Area Applications of ILP to finite element mesh design.

Description of Research The research on ILP at TFM concentrates on the use of ILP methods to build a knowledge base for a finite element mesh design expert system [5], which is one out of three modules that we are proposing for an integrated ES for finite element methods (FEM) [6]. The first experiments have been done in 1990 at the Turing Institute in Glasgow by using the ILP algorithm GOLEM [1]. Respecting the results of the first experiments, a learning set has been slightly changed and extended. The knowledge base was inducted employing GOLEM again [2]. In cooperation with the AI group in Ljubljana a comparison between the results obtained by using three relation learning systems FOIL, GOLEM and LINUS has also been done [3].

The evaluation of the induced knowledge base has shown that the ILP application is an effective approach to solve the finite element mesh design problem. Yet, there are still some steps that have to be taken in order to improve the reliability and transparency of the rules in the knowledge base. Therefore we are planning to continue the research on this project in the directions described in [4]. The major problem in this stage is how to obtain enough appropriate FEM models to be added into the learning set. A wide action all over the Europe has been started recently [7].

Recent Publications

[1] B. Dolšak, S. Muggleton. The application of inductive logic programming to finite element mesh design. In S. Muggleton (ed.) Inductive Logic Programming, 451–472. Academic Press, London, 1992.

[2] B. Dolšak, A. Jezernik, I. Bratko. A knowledge base for finite element mesh design. Proc. International School for the Synthesis of Expert Knowledge, ISSEK Workshop, Bled, 1992.

[3] S. Džeroski, B. Dolšak. Comparison of ILP systems on the problem of finite element mesh design. Proc. International School for the Synthesis of Expert Knowledge, ISSEK Workshop, Bled, 1992.

[4] B. Dolšak, A. Jezernik. Expert system for finite element mesh design. Proc. International Conference Design to Manufacture in Modern Industry, 401–405. DMMI, Bled, 1993.

[5] B. Dolšak, A. Jezernik. A proposal for FEM expert system. Proc. Seventh World Congress on Finite Element Methods, 130–136. Monte Carlo, 1993.

[6] A. Jezernik, B. Dolšak. Expert systems for FEM. Proc. Seventh World Congress on Finite Element Methods, 150–153. Monte Carlo, 1993.

[7] B. Dolšak. Machine learning as a knowledge acquisition method for finite element mesh design. Finite Element News 6, 41–43. Robinson and Associates, Wimborne, Dorset, 1993.

Contact Person Bojan Dolšak Faculty of Technical Sciences Department of Mechanical Engineering Smetanova 17 62000 Maribor Slovenia Phone: +386 62 20091 Fax: +386 62 225451 or 225013 E-mail: dolsak@uni-mb.si